OpenStudy (anonymous):

find the equation of the line useing the following info show all your work (6.3),m-2/3

OpenStudy (anonymous):

is it (6,3)?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

patty got it again...whew!

OpenStudy (anonymous):

(y-ysub1)=m(x-xsub1) (y-3)=-2/3(x-6) Distribute the -2 (y-3)=(-2x+12)/3 3y-9=-2x+12 Do the cross-multiplication 2x+3y=12+9 Transpose -2x and -9 2x+3y=21 This is the equation now. You may check it by substituting (6,3) to this.

OpenStudy (anonymous):

Ms. bio, Thanks for the support.:)